OpenStudy (anonymous):

The half-life of a certain radioactive material is 82 hours. An initial amount of the material has a mass of 166kg. Write an exponential function that models the decay of the material. Find how much radioactive material remains after 12 hours.

OpenStudy (anonymous):

\[166\times (\frac{1}{2})^{\frac{t}{82}}\] should do it

OpenStudy (anonymous):

thank you i didn't realize it was that easy!!